OpenStudy (anonymous):

Let f(x) = 2x+5 and g(x) = x^2-3x+2. Perform each function operation and then find the domain. 3f(x)-2

OpenStudy (anonymous):

is domain is R?

OpenStudy (anonymous):

@kavitha I don't understand what you mean. Domain is usually all real numbers with exceptions

OpenStudy (anonymous):

but here for no value of x this becomes infinity. so is R.

OpenStudy (anonymous):

I guess so. R is real numbers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So all I have to do is, 3f(x)-2=3(2x+5)-2= 6x+15-2 = 6x+13?

OpenStudy (anonymous):

@kavitha

OpenStudy (anonymous):

Is it the one you asking?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

find the domain. 3f(x)-2???

OpenStudy (anonymous):

That' s what it look incompleted because no g(x) involved!

OpenStudy (anonymous):

Look at 48, thats how it is

OpenStudy (anonymous):

I see, it's like random combination question

OpenStudy (anonymous):

so my work is right?

OpenStudy (anonymous):

I need to do it myself: 3f(x) = 3 *(2x+5) = 6x + 15 -> 3f(x)-2 = 6x + 13

OpenStudy (anonymous):

Domain of straight line is R. (line is unlimited)

OpenStudy (anonymous):

I'm going to sleep now, It's almost 4AM my time :(